# LIST: Learning to Index Spatio-Textual Data for Embedding based Spatial Keyword Queries

### 12 Mar 2024 v1

@@ V-D Effect of Dataset Size (RQ3)
(a) The Impact over Effectiveness
(b) The Impact over Training time
Figure 7: The impact of training dataset size over effectiveness and training time on Beijing dataset.
V-E Ablation Studies (RQ4)
Index Study. To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset. We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects. The results are reported in Table VI, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings. We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table IV), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1. (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed. The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table IV.
TABLE VI: Index ablation study on the Geo-Glue dataset.
@@ V-E Ablation Studies (RQ4)
Cluster Quality Study. To illustrate the quality of produced clusters by our index, we conduct a comparison study. We use the proposed relevance model to generate embeddings and then employ our index (denoted as LIST-I) and IVF index to produce clusters separately. We present the cluster results of $IF(C)$ and $P(C)$ in Table VII, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
TABLE VII: Comparison of the quality of clusters.
@@
Pesudo-Label Study. As discussed in Section IV-C, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index. To investigate the impacts of the pseudo-negative labels, we vary the hyperparameter $\text{neg}_{\text{start}}$ from 40,000, 50,000, 60,000, and 70,000, to 80,000 on the Beijing and Shanghai datasets. We illustrate the metrics $P(C)$ and $\text{IF}(C)$ of produced clusters in Figure 8. Notably, as $\text{neg}_{\text{start}}$ increases, both $IF(C)$ and $P(C)$ tend to increase. An increased $IF(C)$ suggests a more concentrated distribution of objects, while an increased $P(C)$ indicates improved accuracy in the retrieval results. The experiment results indicate that the choice of $\text{neg}_{\text{start}}$ leads to a trade-off between effectiveness and efficiency, which can be set flexibly in real-world applications.
@@
[…]
TABLE VIII: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
@@
Weight Learning Study. We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW [5] (denoted as ADrW). Table VIII reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models. Here, LIST-R represents our weight learning mechanism, while LIST-R+ADrW denotes replacing it with the ADrW for training. The results indicate that our weight learning mechanism surpasses the latest ADrW mechanism.
VI CONCLUSIONS AND FUTURE WORK

### 18 Mar 2024 v2

@@ 5.4 Scalibility Study (RQ3)
5.5 Ablation Studies (RQ4)
Index Study. To validate the contribution of our proposed index to the overall effectiveness, we compare it with other baseline indexes, including TkQ, IVF, LSH, and HNSW on the Geo-Glue dataset. We employ each index to retrieve objects and then use our proposed relevance model to rerank these objects. The results are reported in Table 6, where all baselines are denoted as Index+LIST-R and use their default hyperparameter settings. We observe that (1) Compared with the IVF+LIST-R, which has comparative querying speed with LIST (Table 4), LIST obtains a 52.33% improvement on the recall@10 metric and 46.84% improvement on the NDCG@1. (2) Our proposed index consistently outperforms other index baselines and delivers at least an order of magnitude faster querying speed. The runtime of Index+LIST-R is close to the runtime of Index+${}_{D}$, as reported in Table 4.
Table 6: Index ablation study on the Geo-Glue dataset.
@@ 5.5 Ablation Studies (RQ4)
Cluster Quality Study. To illustrate the quality of produced clusters by our index, we conduct a comparison study. We use the proposed relevance model to generate embeddings and then employ our index (denoted as LIST-I) and IVF index to produce clusters separately. We present the cluster results of $IF(C)$ and $P(C)$ in Table 7, which shows that our index achieves much higher precision and obtains comparative imbalance factors compared with IVF index.
Table 7: Comparison of the quality of clusters.
@@
Pesudo-Label Study. As discussed in Section 4.3, the parameter $\text{neg}_{\text{start}}$ affects the difficulty level of pseudo-negative labels, which then impacts the effectiveness and efficiency of our index. To investigate the impacts of the pseudo-negative labels, we vary the hyperparameter $\text{neg}_{\text{start}}$ from 40,000, 50,000, 60,000, and 70,000, to 80,000 on the Beijing and Shanghai datasets. We illustrate the metrics $P(C)$ and $\text{IF}(C)$ of produced clusters in Figure 7. Notably, as $\text{neg}_{\text{start}}$ increases, both $IF(C)$ and $P(C)$ tend to increase. An increased $IF(C)$ suggests a more concentrated distribution of objects, while an increased $P(C)$ indicates improved accuracy in the retrieval results. The experiment results indicate that the choice of $\text{neg}_{\text{start}}$ leads to a trade-off between effectiveness and efficiency, which can be set flexibly in real-world applications.
@@
Spatial Learning Study. To evaluate the learning-based spatial relevance module, we consider the following baseline: (1) LIST-R+$S_{in}$ that replaces the learning-based spatial relevance module with $S_{in}$ for training, where $S_{in}$ = 1 - $SDist(q.loc,o.loc)$; and (2) LIST-R+$\alpha*S_{in}^{\beta}$ which substitutes the spatial relevance module with a learnable exponential function. $\alpha$ and $\beta$ are two learnable parameters and are processed to ensure non-negative. Table 8 presents the experimental results obtained by conducting a brute-force search on the Beijing dataset using the trained models. Notably, LIST-R outperforms LIST-R+$S_{in}$ in all metrics. Interestingly, the first variant outperforms the second variant, which suggests that without careful design, a learnable function may be outperformed by a simple approach.
Table 8: Ablation study of spatial relevance module and weight learning via brute-force search on the Beijing dataset.
@@
Weight Learning Study. We conduct a comparison experiment between our weight learning module and the attention mechanism proposed by DrW liu2023effectiveness (denoted as ADrW). Table 8 reports the results obtained by conducting a brute-force search on the Beijing dataset using the trained relevance models. Here, LIST-R represents our weight learning mechanism, while LIST-R+ADrW denotes replacing it with the ADrW for training. The results indicate that our weight learning mechanism surpasses the latest ADrW mechanism.
(a) The Impact over Effectiveness
(b) The Impact over Training time
Figure 8: The impact of training dataset size over effectiveness and training time on the Beijing dataset.
6 Conclusions and Future Work